# Can GenAI Move from Individual Use to Collaborative Work? Experiences, Challenges, and Opportunities of Integrating GenAI into Collaborative Newsroom Routines

### v2

@@ 8.3. Design and Policy Implication
We call for bringing GenAI use into newsrooms out of the shadows and embedding it into collaborative, transparent, and accountable practices that align with both everyday workflows and institutional values.
@@ 8.3.1. Designing for Collaborative GenAI Integration
First, design for social traceability to counter the “double black box.” In our study, GenAI use was often invisible on two levels: editors and colleagues rarely knew when or how journalists were using GenAI, and journalists themselves often lacked clarity on how the model produced its outputs. This double opacity not only undermines accountability but also weakens trust in collaborative processes. Design should therefore make both human and AI contributions more visible across the editorial workflow. For instance, interfaces could visually differentiate between human-generated and AI-generated text, display prompt histories, and show revision paths.
Second, design for shared editorial work flow. Current GenAI tools primarily operate in a single-user paradigm, where one journalist privately interacts with the system. Yet our participants increasingly envisioned GenAI as a shared infrastructure woven into team-based routines. To realize this vision, design should enable collaborative use: shared GenAI workspaces, prompt libraries editable by multiple users, and co-edit outputs that record collective contributions.
Third, turning institutional norms into interface-level affordance. Journalists sought tools that would help them uphold values of transparency, attribution, and editorial integrity. Systems can support this by prompting disclosure of GenAI involvement, suggesting attribution formats consistent with newsroom policies, or flagging content that might raise ethical concerns. Integrating these practices into the interface itself would allow journalists to enact shared values in everyday work, helping normalize GenAI as a legitimate, collectively governed part of newsroom routines.
@@ 8.3.2. Toward Collaborative and Accountable GenAI Policies
First, to meaningfully support GenAI-augmented collaboration, newsrooms must treat collaboration not as a downstream usability concern, but as a strategic design priority. Institutional policies should incentivize the early involvement of front-line journalists in AI development processes and require that GenAI evaluation metrics extend beyond technical benchmarks like accuracy. Instead, assessments should include collaborative usability, editorial interpretability, and shared accountability across team members.
Second, newsrooms must establish clearer norms and incentives to institutionalize the responsible use of GenAI. Our findings indicate that informal, often stigmatized use of GenAI persists in part because of the absence of formal guidance and supportive peer practices. Policies should not treat GenAI use as a compliance problem, but as a cultural opportunity to scaffold shared learning and responsible exploration. Practices such as shared prompt libraries, peer-learning forums, and transparent audit trails can foster a culture of open dialogue, where GenAI use is not hidden but collaboratively refined and improved.
Third, for GenAI to be governable and sustainable in the long term, it must be integrated into the newsroom’s editorial routines. We emphasize that GenAI adoption and integration must occur at the infrastructural level. Policies should require traceable logging of GenAI involvement in content production, guidelines for data retention, and clear delineation of editorial responsibility when AI-generated content is used. These mechanisms not only support internal accountability and collaboration, but also prepare news organizations to meet external demands, such as regulatory compliance, reputational risk management, and public scrutiny over AI usage in high-stakes reporting environments.
@@ 9. Limitations and Future Work
This study provides early insight into the emerging role of GenAI in newswork collaboration, but several limitations should be noted. First, our sample is limited to Chinese newsrooms, which operate within a distinct media system shaped by strong institutional norms and political constraints. While many of the structural and cultural dynamics we observed may resonate globally, future research should examine how GenAI adoption varies across different media ecosystems, including commercial, public service, and nonprofit news organizations in democratic contexts. Examining related sectors such as education, law, or healthcare could also offer valuable contrasts in how GenAI is adopted into collaborative practices in high-stakes domains.
Second, our recruitment strategy and study design may have favored journalists who used GenAI directly and were relatively open or positive toward GenAI. Because many participants were recruited through personal networks and snowball sampling, those who were skeptical, disengaged, or more deeply concerned about GenAI’s risks may be underrepresented. This limitation is compounded by the reliance on self-reported accounts, which can introduce biases in how participants frame their own practices and attitudes. Future work should seek to incorporate more diverse perspectives, including journalists who resist or avoid GenAI, and combine self-reports with direct observation or usage data to mitigate these biases.
Finally, our research focused primarily on newsroom staff, particularly editors and journalists. However, collaborative GenAI use may also involve non-editorial actors such as designers, developers, and product managers. Future work should adopt a more holistic organizational lens to trace how various teams negotiate GenAI integration.
@@ 10. Conclusion
Through 27 interviews with senior editors, newsroom managers and front-line journalists in China, this study examined how journalists navigate the integration of GenAI into collaborative newswork. We identified a clear gap between individual experimentation with GenAI and the structural and cultural mechanisms needed to support sustained, legitimate, and team-based use. While journalists expressed curiosity and even enthusiasm about GenAI’s potential, their current practices remain fragmented, private, and stigmatized. Our paper contributes to HCI and CSCW by foregrounding the sociotechnical barriers that prevent GenAI from becoming a collaborative infrastructure and pointing out future opportunities to integrate GenAI into collaborative work.

### v3

@@ 8.4. Design and Policy Implication
We call for bringing GenAI use into newsrooms out of the shadows and embedding it into collaborative, transparent, and accountable practices that align with both everyday workflows and institutional values. Building on our findings, we argue that better coordinating GenAI into collaboration in newsrooms requires interventions at two levels: (1) bottom-up design that equips everyday users with tools for transparency, coordination, and shared control, and (2) top-down policy structures that institutionalize accountability and align GenAI use with organizational values. This dual perspective reflects longstanding HCI guidance that successful human–AI systems could integrate interaction-level and institution-level interventions. Table 2 summarizes the key implications that we elaborate in the following two subsections, distinguishing bottom-up design considerations from top-down policy needs.
Table 2. Summary of Key Design and Policy Implications for Collaborative GenAI Coordination in Newsrooms
Key Challenge Area Bottom-Up Design Implications Top-Down Policy Implications
Transparency and Traceability (marked by 6.1.2, 6.2, 7.2) Make human–AI contributions visible across workflows; provide prompt histories, revision paths, and provenance indicators. Require traceable logging of GenAI involvement, retention policies, and disclosure guidelines aligned with editorial standards.
Collaborative Workflows (marked by 6.1.1) Design shared GenAI workspaces, multi-user prompt libraries, and co-editing environments that reflect newsroom teamwork. Define organizational expectations for collaborative GenAI use; incentivize early front-line journalist participation in AI tool development and evaluation.
Institutional Norms as Interface Affordances (marked by 6.2) Embed newsroom values into interface-level prompts: e.g., reminders for attribution, accuracy checks, and ethical flags. Establish formal norms for responsible GenAI use; build institutional structures for transparency, auditing, and consistent editorial review.
Learning and Trust-Building (marked by 5.2, 6.2) Support gradual user learning with features enabling peer knowledge sharing and cross-role visibility. Create shared prompt libraries, peer-learning forums, and organizational training programs supporting GenAI literacy.
Accountability Structures (marked by 6.1.2, 6.1.3) Design mechanisms to clarify responsibility for AI-assisted content, especially in multi-party editing tasks. Develop clear accountability frameworks specifying editorial responsibility, escalation paths, and how AI errors are handled.
Scalability and Governance (marked by 6.1.2) Build interaction-level affordances that scale across teams (e.g., configurable transparency levels for editors vs. journalists). Integrate GenAI into editorial infrastructures; prepare for regulatory demands and reputational risk management in high-stakes reporting.
@@ 8.4.1. Designing for Collaborative GenAI Integration
From a bottom-up perspective, effective collaborative GenAI coordination could begin with interaction-level design that supports journalists’ real practices.
First, design for social traceability to counter the “double black box.” In our study, GenAI use was often invisible on two levels: editors and colleagues rarely knew when or how journalists were using GenAI, and journalists themselves often lacked clarity on how the model produced its outputs. This double opacity not only undermines accountability but also weakens trust in collaborative processes. Design could therefore make both human and AI contributions more visible across the editorial workflow. For instance, interfaces could visually differentiate between human-generated and AI-generated text, display prompt histories, and show revision paths. This design stance aligns with Human–AI Interaction guidelines that call for visible system reasoning and meaningful traceability by translating them into concrete mechanisms for editorial collaboration, clarifying how human-AI actions interact with human judgment across the workflow (Amershi et al., 2019).
Second, design for shared editorial workflow. Current GenAI tools primarily operate in a single-user paradigm, where one journalist privately interacts with the system. Yet our participants increasingly envisioned GenAI as a shared infrastructure woven into teamwork. To realize this vision, design could enable collaborative use: shared GenAI workspaces, prompt libraries editable by multiple users, and co-edit outputs that record collective contributions. These ideas resonate with existing HCI design recommendations encouraging systems to support collaborative use cases and enable cross-role coordination in AI-augmented work (Amershi et al., 2019; Deng et al., 2023; Xiao et al., 2025c).
Third, turning institutional norms into interface-level affordances. Journalists sought tools that would help them uphold values of transparency, attribution, and editorial integrity. Systems can support this by prompting disclosure of GenAI involvement, suggesting attribution formats consistent with newsroom policies, or flagging content that might raise ethical concerns. Integrating these practices into the interface itself would allow journalists to enact shared values in everyday work, helping normalize GenAI as a legitimate, collectively governed part of newswork. This design direction follows Human–AI guidelines such as “match relevant social norms” and ”mitigate social biases” which are especially important in high-stakes domains like journalism (Li et al., 2025c; Amershi et al., 2019).
@@ 8.4.2. Toward Collaborative and Accountable GenAI Policies
While bottom-up design interventions can scaffold everyday collaborative use, they could be complemented by top-down organizational policies that institutionalize responsible GenAI use at scale. Without structural support, interaction-level transparency or coordination mechanisms cannot fully address issues of accountability, risk, and uneven power dynamics identified in our findings.
First, to meaningfully support coordination of GenAI into collaborative newswork, newsrooms could treat collaboration not as a downstream usability concern, but as a strategic design priority. Institutional policies could incentivize the early involvement of front-line journalists in AI development processes and require that GenAI evaluation metrics extend beyond technical benchmarks like accuracy. Instead, assessments could include collaborative usability, editorial interpretability, and shared accountability across team members. Such alignment between organizational policies and interface-level design mirrors the Human–AI Interaction guideline encouraging alignment between AI capabilities and organizational expectations (Li et al., 2025b; Murire, 2024; Rakova et al., 2021).
Second, newsrooms could establish clearer norms and incentives to institutionalize the responsible use of GenAI. Our findings indicate that informal, often stigmatized use of GenAI persists in part because of the absence of formal guidance and supportive peer practices. Policies could not treat GenAI use as a compliance problem, but as a cultural opportunity to scaffold shared learning and responsible exploration. Practices such as shared prompt libraries, peer-learning forums, and transparent audit trails can foster a culture of open dialogue, where GenAI use is not hidden but collaboratively refined and improved. This echoes the HAI guideline to support user learning across time, which emphasizes that GenAI competence could grow within a social and organizational context (Amershi et al., 2019; Murire, 2024; Li et al., 2025b).
Third, for GenAI to be governable and sustainable in the long term, it could be integrated into the newsroom’s editorial newswork. We emphasize that GenAI adoption and coordination could occur at the infrastructural level. Policies could require traceable logging of GenAI involvement in content production, guidelines for data retention, and clear delineation of editorial responsibility when AI-generated content is used. These mechanisms not only support internal accountability and collaboration, but also prepare news organizations to meet external demands, such as regulatory compliance, reputational risk management, and public scrutiny over AI usage in high-stakes reporting environments. By pairing these top-down governance frameworks with bottom-up design changes, newsrooms can learn from HCI scholars and move toward a socio-technical ecosystem that supports transparent, accountable, and trustworthy GenAI-enabled collaboration (Aitamurto et al., 2019).
@@ 9. Limitations and Future Work
This study provides early and exploratory insight into the emerging role of GenAI in collaborative newswork. As an exploratory qualitative investigation, our findings illuminate emergent patterns in how journalists personally experiment with GenAI and how these practices create both tensions and opportunities for collaboration. However, this exploratory nature also brings several limitations that point toward important directions for future research.
First, our sample is limited to Chinese newsrooms, which operate within a distinct media system shaped by strong institutional norms and political constraints. While many of the structural and cultural dynamics we observed may resonate globally, our findings could not be interpreted as universally representative. Future work could examine how GenAI adoption varies across different media ecosystems, including commercial, public service, and nonprofit news organizations in democratic contexts. Comparative studies could further explore how political systems, regulatory environments, and organizational cultures shape the institutionalization of GenAI-supported collaboration. Examining related sectors such as education, law, or healthcare may also reveal contrasting pathways for incorporating GenAI into high-stakes collaborative practices.
Second, our recruitment strategy and study design may have favored journalists who were already using GenAI directly and were relatively open or positive toward experimenting with it. Because many participants were recruited through personal networks and snowball sampling, those who are skeptical, disengaged, or more deeply concerned about GenAI may be underrepresented. This limitation is compounded by the reliance on self-reported accounts, which can introduce biases in how participants frame their own practices and attitudes. Future work could seek to incorporate more diverse perspectives, including journalists who resist, avoid, or strategically withhold GenAI use, and could combine interviews with direct observation, trace data, or system logs to mitigate self-report bias.
Third, while our findings speculate on possible future workflows and organizational coordination, we lack deployment-based or behavioral data to demonstrate whether the proposed collaborative mechanisms would indeed enhance multi-party trust, coordination, or efficiency in practice. Our analysis identifies potential workflows and cultural transformations, but empirical evidence from real-world deployment or longitudinal observation is needed to validate whether these workflows truly support collaborative accountability, reduce invisible labor, or shift newsroom culture. Future work could therefore explore the design, prototyping, and in situ deployment of collaborative GenAI systems, studying how such systems are appropriated, resisted, or reshaped over time.
Finally, although our study focused primarily on editorial staff, particularly journalists and editors, collaborative GenAI use in news organizations may extend beyond editorial roles. Designers, developers, product managers, and data teams increasingly contribute to the integration of GenAI in newsroom infrastructures. Future work could adopt a more holistic organizational lens to trace how cross-functional teams negotiate GenAI integration, how institutional priorities shape design and deployment decisions, and how organizational boundaries and power asymmetries influence collaborative uptake.
@@ 10. Conclusion
Through 27 interviews with senior editors, newsroom managers and front-line journalists in China, this study examined how journalists navigate the coordination of GenAI into collaborative newswork. We identified a clear gap between individual experimentation with GenAI and the structural and cultural mechanisms needed to support sustained, legitimate, and team-based use. While journalists expressed curiosity and even enthusiasm about GenAI’s potential, their current practices remain fragmented, private, and stigmatized. Our paper contributes to HCI and CSCW by foregrounding the sociotechnical barriers that prevent GenAI from becoming a collaborative infrastructure and pointing out future opportunities to coordinate GenAI into collaborative work.
Appendix A Interview Protocol and Demographic Questions
This appendix provides the full interview protocol and the demographic questions used to contextualize participants’ backgrounds and collaborative roles. The protocol followed a semi-structured, three-phase design, allowing researchers to probe deeply into participants’ concrete practices, expectations, and critical reflections on GenAI in collaborative newswork.
A.1. Demographic Questions
At the beginning of each interview, participants were asked a brief set of demographic questions to contextualize their professional background and position within the newsroom. These questions were optional and used solely to support analysis of GenAI practices across different roles and experience levels.
  • Gender: How do you identify your gender?
  • Age: What is your age?
  • Years in Journalism: How many years have you worked in journalism or media production?
  • Current Position: What is your role in the newsroom?
  • Newsroom Unit or Department: Which desk, team, or section are you part of? (e.g., society, politics, investigations, video)
  • Primary Collaborative Partners: With whom do you most often collaborate in daily newswork?
  • Experience with GenAI Tools:
    • Have you used GenAI tools in your reporting or editing work?
    • If yes, for which tasks?
    • Frequency of use: (Several times per day / Daily / Several times per week / Occasionally)
  • Involvement in GenAI Decision-Making:
    • Are you involved in discussions or planning related to GenAI adoption in your newsroom?
A.2. Interview Protocol
The interviews were semi-structured and organized into three main phases. Follow-up probing questions were used as appropriate to deepen or clarify participants’ accounts.
A.2.1. Phase 1: Concrete Experiences with GenAI in Newswork
  • Can you describe specific moments when you used GenAI in your daily reporting or editing work?
  • What types of GenAI tools or platforms did you use, and how did you access them (e.g., internal tools, public LLMs)?
  • At which points of the reporting or editing process did you most commonly use GenAI (e.g., idea generation, background research, drafting, editing, verification)?
  • How did GenAI use unfold in collaborative contexts, for example with editors or teammates?
  • Were there moments when you used GenAI together with colleagues, or shared AI-generated content with others? How did they respond?
  • Can you recall situations where GenAI changed how responsibility or workload was distributed within a team?
  • In what ways did GenAI fit or fail to fit into your existing newsroom routines?
  • Did GenAI use ever create uncertainty or confusion in collaborative work (e.g., accountability, authorship, reliability)?
  • How did your organization communicate or coordinate early-stage GenAI practices? Were there any informal norms or expectations?
  • How did GenAI shape your interactions with editors, colleagues, or other teams?
  • Have you ever chosen not to use GenAI in a situation where you could have? Why?
  • Have you witnessed colleagues using GenAI in ways that affected your own work—positively or negatively?
  • How did GenAI affect the pace, accuracy, or style of your reporting or editing tasks?
  • Did GenAI introduce any new forms of invisible work (e.g., checking, correcting, validating outputs)?
A.2.2. Phase 2: Expectations for Collaborative GenAI Integration
  • In your ideal scenario, how would GenAI support your team’s collaborative work?
  • What kinds of assistance would you want GenAI to provide for teamwork or cross-role coordination?
  • What would make a GenAI tool feel reliable, trustworthy, or compatible with collaborative newswork?
  • How would you envision GenAI interacting with editors, senior managers, or fellow journalists?
  • If GenAI could understand your team’s workflow, what aspects could it pay attention to?
  • What would an “ideal” GenAI system know about your newsroom’s editorial standards or organizational values?
  • How proactively could GenAI participate in teamwork?
  • What kinds of transparency would you need from GenAI to feel comfortable using it collaboratively?
  • How could GenAI tools display provenance, e.g., where ideas, edits, or facts came from, when supporting collaborative tasks?
  • Do you imagine different roles in the newsroom using GenAI differently? How could tools adapt to these differences?
  • What would an optimal division of labor between humans and GenAI look like in reporting, editing, and publishing?
  • If you could redesign your current workflow with GenAI in mind, what changes would you make?
  • How could GenAI support cross-team workflows?
  • Are there any risks or unintended consequences that an ideal collaborative GenAI system could help prevent?
  • What kinds of training or onboarding would help your team use GenAI more effectively together?
  • How could GenAI tools negotiate ethical concerns?
  • In what ways could GenAI strengthen teamwork culture, e.g., coordination, communication, shared understanding?
A.2.3. Phase 3: Gaps Between Ideals and Current Practices
  • Why do you think GenAI has not yet become a regular part of your team’s workflow?
  • What challenges, breakdowns, or tensions arise when colleagues use—or do not use—GenAI?
  • What forms of uncertainty, hesitation, or resistance have you observed in your team?
  • What organizational or technological changes would make GenAI more compatible with collaborative newsroom practices?
  • What concerns or risks do you foresee in deeper GenAI integration?
  • Can you recall any specific instances when GenAI created confusion, misalignment, or extra coordination work?
  • Have you experienced situations where teammates used GenAI but did not disclose it? How did this affect collaboration?
  • What makes it difficult for colleagues to trust GenAI—or to trust each other when using GenAI?
  • How do power dynamics (e.g., junior journalists vs. editor) affect willingness to use or share GenAI-generated content?
  • How does the lack of clear guidelines or protocols influence GenAI use in your team?
  • Do you think the current newsroom workflow or CMS makes it hard to incorporate GenAI into collaboration?
  • Are there concerns about accountability when GenAI is used in reporting or editing?
  • What kinds of errors or inaccuracies have you seen from GenAI that created extra work for your team?
  • Are there legal, political, or reputational risks that discourage GenAI adoption in your newsroom?
  • Are there differences between how willing different roles (e.g., editors, junior journalists, managers) are to use GenAI?
  • Have organizational priorities or incentives made it harder (or easier) to integrate GenAI into team workflows?
  • What prevents GenAI from supporting more complex collaborative tasks?
  • If you imagine your ideal workflow again, what are the biggest gaps between that vision and your current reality?
Notes
  • The protocol was used flexibly and adapted to participants with different newsroom roles.